OpenStudy (anonymous):

Find the set of points that are part of a line that is perpendicular to the line y = ­-3x

jimthompson5910 (jim_thompson5910):

hint: original slope = a/b perpendicular slope = -b/a notice how I flipped the fraction and the sign

jimthompson5910 (jim_thompson5910):

example original slope: 2/7 perpendicular slope: -7/2

OpenStudy (jdoe0001):

what's the slope of \(\bf y = ­-3x?\)

OpenStudy (anonymous):

@jim_thompson5910 Ah yeah so the reciprocal?

jimthompson5910 (jim_thompson5910):

the negative reciprocal

OpenStudy (anonymous):

Right, I see. Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome